Circular No. 5957

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA CASSIOPEIAE 1993 C. Mason and R. D. Gehrz, University of Minnesota, write: “We report

infrared magnitudes for N Cas 1993 obtained on Mar. 22.9 UT with a bolometer on the 0.76-m telescope at the O’Brien Observatory using a 27′′ beam and a 34′′ throw: K = 4.6 ± 0.16, L = 1.23 ± 0.02, M = −0.06 ± 0.01, [8.7 μm] = −1.16 ± 0.19, [9.8 μm] = −1.28 ± 0.17, [10.3 μm] = −0.82 ± 0.31. The measurements conform to a 690-K blackbody and show that dust is present in the ejecta. The relatively low grain temperature suggests that the dust condensed some time ago.”

Further photoelectric photometry by H. Mikuž, Ljubljana, Slovenia (cf. IAUC 5945): Mar. 11.78 UT, V = 15.99 ± 0.05.

(2060) CHIRON M. P. Womack, Northern Arizona University; and S. A. Stern, South-

west Research Institute, report: “We have obtained millimeter-wave observations resulting in upper limits on the CO column abundance in the coma of (2060) Chiron. These observations were made during Feb. 25–26 UT using the Caltech Submillimeter Observatory on Mauna Kea; we observed the CO J = 2–1 transition at a frequency of 230 GHz (1.3 mm). Data were obtained with resolutions of both 50 and 500 kHz in good observing conditions (i.e., τ ∼ 0.08). We achieved 3σ upper limits for the 2–1 emission of 75 and 35 mK at the 50- and 500-kHz resolutions, respectively. Assuming (i) Chiron’s coma fills the 30′′ CSO beam and (ii) the gas excitation, rotational, and kinetic temperatures are all between 10 and 50 K, then for both resolutions, the upper limit corresponds to an average CO column density in the beam of < (3 ± 1) × 10^{13} cm^{−2}. A Haser model calculation, assuming an isotropic coma expansion at velocities of 0.2–0.5 km s^{−1} and a CO line-width twice the coma expansion velocity, implies production-rate upper limits of Q_{CO} < 6.5–15 × 10^{27} s^{−1}, respectively, if CO is a parent molecule, and < 9.0–36 × 10^{27} s^{−1} if CO is the daughter product of an H_{2}CO parent. If the coma is not expanding isotropically, and instead the CO J = 2–1 line widths are much narrower (e.g., ∼ 1/3 the coma expansion velocity), then the derived column densities and CO production limits could be up to five times lower.”

1994 March 25 Daniel W. E. Green